OpenStudy (anonymous):

If the parent function is y=1/x. describe the change in the equation y= 1/(x+4)

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

Its multiple choice too =)

OpenStudy (anonymous):

y = 1/(x+4) looks exactly the same as y = 1/x except that it is shifted 4 units to the left.

OpenStudy (anonymous):